# Hyper-Transforming Latent Diffusion Models

### v3

@@ 1 Introduction
×$\times 0.125$
×$\times 0.25$
×$\times 0.5$
×$\times 1$
×$\times 2$
×$\times 4$
@@ 2.1 Latent Diffusion Models
[…]
@@ 3.1 Notation
We define =$\mathcal{F}=\{f:\mathcal{X}\to\mathcal{Y}\}$ as the space of continuous signals, where $\mathcal{X}$ represents the domain of coordinates, and $\mathcal{Y}$ the codomain of signal values, or features. Let $\mathcal{D}$ be a dataset consisting of $N$ pairs $\mathcal{D}\doteq\{({\bm{X}}_{n},{\bm{Y}}_{n})\}_{n=1}^{N}$. Here, ${\bm{X}}_{n}$ and ${\bm{Y}}_{n}$ conform a signal as a collection of $D_{n}$ coordinates ${\bm{X}}_{n}\doteq\{{\bm{x}}^{(n)}_{i}\}_{i=1}^{D_{n}},\;{\bm{x}}^{(n)}_{i}\in\mathcal{X}$ and signal values ${\bm{Y}}_{n}\doteq\{{\bm{y}}^{(n)}_{i}\}_{i=1}^{D_{n}},\;{\bm{y}}^{(n)}_{i}\in\mathcal{Y}$.
@@ 3.2.1 Hyper-Transformer Decoder
${\bm{w}}_{c}=\mathcal{R}({\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}},\bar{{\bm{w}}}_{c}^{\text{b}})$ (7)
@@
$\mathcal{R}^{\text{(norm)}}\left({\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}},\bar{{\bm{w}}}_{c}^{\text{b}}\right)=\frac{{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}\odot\bar{{\bm{w}}}_{c}^{\text{b}}}{\left\|{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}\odot\bar{{\bm{w}}}_{c}^{\text{b}}\right\|}.$ (8)
However, we found this approach unsuitable for generating weights in INRs with periodic activations. We hypothesize that the normalization in Equation (8) causes the resulting weights and their gradients to vanish when $||{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}||\approx 0$, leading to training instability.
@@
$\mathcal{R}^{(\text{scale})}\left({\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}},\bar{{\bm{w}}}_{c}^{\text{b}}\right)=(1+{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}})\odot\bar{{\bm{w}}}_{c}^{\text{b}}.$ (9)
This formulation ensures that when $||{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}||\approx 0$, the base weights remain unchanged, preventing instability in the generated INR parameters.
It is essential to distinguish between the two sets of globally shared learnable parameters. The sequence $\bar{{\mathbf{W}}}_{1:L}^{\text{i}}$ (of length $G\times L$) serves as the initialization for the grouped weight tokens and is provided as input to the Transformer Decoder. In contrast, the sequence $\bar{{\mathbf{W}}}_{1:L}^{\text{b}}$ (of length $d_{in}\times L$) acts as a global reference for the full set of INR weights and is used exclusively for the reconstruction in Equation (9).
@@ 3.4 Latent Diffusion
Figure 5: Uncurated samples from GASP, Functa (diffusion-based) and our LDMI trained with CelebA-HQ at 64$\times$64 resolution (a). In (b), our model was trained on CelebA-HQ at 256$\times$256 using the hyper-transforming approach.
@@ 4.2 Generation
In Figure 5, we provide samples from LDMI trained on CelebA at 64$\times$64 and 256$\times$256, demonstrating its ability to generate high-quality and diverse images across different scales, and the qualitative superiority against the baselines. Additionally, in Table 1, we report FID scores of our samples for CelebA and ImageNet, highlighting the model’s performance in terms of image quality and reconstruction accuracy. Importantly, while GASP achieves a lower FID score on CelebA-HQ, it cannot perform reconstructions due to its adversarial design. In contrast, our LDMI supports both high-quality sampling and accurate reconstructions, as demonstrated in the following experiment.
@@ 4.3 Reconstruction
As for generation, our model enables seamless reconstruction at arbitrary resolutions. Figures 9 and 10 show results from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. To achieve this, we encode the observations, sample from the posterior distribution, and transform these latent samples into INRs using our HD decoder. The resulting INRs are then evaluated on grids of varying resolutions to produce reconstructions ranging from ×$\times 1/8$ to ×$\times 4$. Importantly, such flexible resolution control is not supported by methods like GASP, which cannot perform reconstructions due to their adversarial design.
To further validate our claims regarding cross-modality generalization, we evaluate LDMI on two additional diverse datasets: ShapeNet Chairs and ERA5 climate data. As shown in Table 2, LDMI achieves higher reconstruction quality across both datasets, outperforming Functa by 9.6 dB on Chairs and 9.7 dB on ERA5, while also improving upon VAMoH by 0.4 dB and 5.6 dB respectively. The substantial performance gain on the ERA5 climate dataset (44.6 dB vs. 39.0 dB for VAMoH) highlights LDMI’s exceptional capability in representing complex, continuous spatio-temporal signals beyond standard visual data, further supporting our argument for a truly general-purpose INR-based generative framework.
Model Chairs (dB) $\uparrow$ ERA5 (dB) $\uparrow$
Functa (Dupont et al., 2022a) 29.2 34.9
VAMoH (Koyuncu et al., 2023) 38.4 39.0
LDMI 38.8 44.6
Table 2: Reconstruction quality (PSNR in dB) on ShapeNet Chairs and ERA5 climate data, demonstrating LDMI’s strong generalization capabilities across modalities. Note that GASP is omitted as it is not applicable to INR reconstruction tasks.
@@ Acknowledgements
This research was supported by the Villum Foundation through the Synergy project number 50091, and by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (MLLS, grant NNF20OC0062606).
Ignacio Peis acknowledges support by Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429). Batuhan Koyuncu is supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
@@ A.2 Diffusion Models
In DDPM, the forward diffusion process incrementally corrupts the observed variable ${\bm{z}}_{0}$ using a Markovian noise process with latent variables ${\bm{z}}_{1:T}$
$q({\bm{z}}_{1:T}|{\bm{z}}_{0}):=\prod_{t=1}^{T}q({\bm{z}}_{t}|{\bm{z}}_{t-1}),$ (18)
where $q({\bm{z}}_{t}|{\bm{z}}_{t-1}):=\mathcal{N}({\bm{z}}_{t};\sqrt{1-\beta_{t}}{\bm{z}}_{t-1},\beta_{t}\mathbf{I})$, and $\beta_{t}$ is a noise schedule controlling the variance at each time step.
@@
$p_{\theta}({\bm{z}}_{0:T}):=p({\bm{z}}_{T})\prod_{t=1}^{T}p_{\theta}({\bm{z}}_{t-1}|{\bm{z}}_{t})$ (19)
where $p_{\theta}({\bm{z}}_{t-1}|{\bm{z}}_{t}):=\mathcal{N}\left(\mu_{\theta}({\bm{z}}_{t},t),\bm{\sigma}_{t}^{2}\right)$ is defined as a Gaussian whose mean
@@ C.2 Samples at Multiple Resolutions
Figure 8 displays generations from LDMI trained on data across different modalities, rendered at scales ranging from ×$\times 0.125$ to ×$\times 4$. These results highlight the resolution-agnostic nature of our approach and its ability to produce coherent, high-quality outputs across a wide range of spatial resolutions.
@@
3D Chairs Refer to caption ×$\times 0.125$ Refer to caption ×$\times 0.25$ Refer to caption ×$\times 0.5$ Refer to caption ×$\times 1$ Refer to caption ×$\times 2$ Refer to caption ×$\times 4$
@@ C.3 Reconstructions at Multiple Resolutions
Figures 9 and 10 present reconstructions from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. Reconstructions are rendered at resolutions ranging from ×$\times 0.125$ to ×$\times 4$, showcasing the model’s ability to maintain spatial coherence and fine details even under extreme upsampling. Notably, such flexible reconstruction is not supported by baseline methods like GASP due to architectural limitations. Furthermore, datasets of this complexity are not even addressed by other reconstruction-capable baselines such as VAMoH (Koyuncu et al., 2023), Functa (Dupont et al., 2022b), or Spatial Functa (Bauer et al., 2023).
@@
×$\times 0.125$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 0.25$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 0.5$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 1$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 2$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 4$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
×$\times 0.125$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 0.25$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 0.5$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 1$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 2$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
×$\times 4$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix D Experimental Details
resolution $64\times 64$ $64\times 64$ $256\times 256$ $256\times 256$ $46\times 90$ $32\times 32\times 32$
@@
latent_size $16\times 16$ $16\times 16$ $64\times 64$ $64\times 64$ $11\times 22$ $4\times 4$

### v4

@@ 1 Introduction
$\times 0.125$
$\times 0.25$
$\times 0.5$
$\times 1$
$\times 2$
$\times 4$
@@ 2.1 Latent Diffusion Models
In the second stage, LDMs approximate the aggregate posterior $q_{\psi}({\bm{z}})=\int q_{\psi}({\bm{z}}|{\bm{y}})p_{\text{data}}({\bm{y}})\operatorname{d\!}{\bm{y}}$ with a diffusion model $p_{\theta}({\bm{z}})$, shifting the generative modeling task from data space to latent space and significantly reducing computational cost. While both continuous-time SDE-based approaches (Song et al., 2021b) and discrete Markov chain formulations (Ho et al., 2020; Song et al., 2021a) are viable, we follow the DDPM-based approach as adopted in the original LDM implementation (Rombach et al., 2022).
@@ 3.1 Notation
We define $\mathcal{F}=\{f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{Y}\}$ as the space of continuous signals, where $\mathcal{X}$ represents the domain of coordinates, and $\mathcal{Y}$ the codomain of signal values, or features. Let $\mathcal{D}$ be a dataset consisting of $N$ pairs $\mathcal{D}\doteq\{({\bm{X}}_{n},{\bm{Y}}_{n})\}_{n=1}^{N}$. Here, ${\bm{X}}_{n}$ and ${\bm{Y}}_{n}$ conform a signal as a collection of $D_{n}$ coordinates ${\bm{X}}_{n}\doteq\{{\bm{x}}^{(n)}_{i}\}_{i=1}^{D_{n}},\;{\bm{x}}^{(n)}_{i}\in\mathcal{X}$ and signal values ${\bm{Y}}_{n}\doteq\{{\bm{y}}^{(n)}_{i}\}_{i=1}^{D_{n}},\;{\bm{y}}^{(n)}_{i}\in\mathcal{Y}$.
@@ 3.2.1 Hyper-Transformer Decoder
${\bm{w}}_{c}=\mathcal{R}({\bm{w}}_{\lceil c/k\rceil}^{\text{o}},\bar{{\bm{w}}}_{c}^{\text{b}})$ (7)
@@
$\mathcal{R}^{\text{(norm)}}\left({\bm{w}}_{\lceil c/k\rceil}^{\text{o}},\bar{{\bm{w}}}_{c}^{\text{b}}\right)=\frac{{\bm{w}}_{\lceil c/k\rceil}^{\text{o}}\odot\bar{{\bm{w}}}_{c}^{\text{b}}}{\left\|{\bm{w}}_{\lceil c/k\rceil}^{\text{o}}\odot\bar{{\bm{w}}}_{c}^{\text{b}}\right\|}.$ (8)
However, we found this approach unsuitable for generating weights in INRs with periodic activations. We hypothesize that the normalization in Equation (8) causes the resulting weights and their gradients to vanish when ${\bm{w}}_{\lceil c/k\rceil}^{\text{o}}||\approx 0$, leading to training instability.
@@
$\mathcal{R}^{(\text{scale})}\left({\bm{w}}_{\lceil c/k\rceil}^{\text{o}},\bar{{\bm{w}}}_{c}^{\text{b}}\right)=(1+{\bm{w}}_{\lceil c/k\rceil}^{\text{o}})\odot\bar{{\bm{w}}}_{c}^{\text{b}}.$ (9)
This formulation ensures that when ${\bm{w}}_{\lceil c/k\rceil}^{\text{o}}||\approx 0$, the base weights remain unchanged, preventing instability in the generated INR parameters.
It is essential to distinguish between the two sets of globally shared learnable parameters. The sequence $\bar{{\mathbf{W}}}_{1\mathrel{\mathop{\ordinarycolon}}L}^{\text{i}}$ (of length $G\times L$) serves as the initialization for the grouped weight tokens and is provided as input to the Transformer Decoder. In contrast, the sequence $\bar{{\mathbf{W}}}_{1\mathrel{\mathop{\ordinarycolon}}L}^{\text{b}}$ (of length $d_{in}\times L$) acts as a global reference for the full set of INR weights and is used exclusively for the reconstruction in Equation (9).
@@ 3.4 Latent Diffusion
Figure 5: Uncurated samples from GASP, Functa (diffusion-based) and our LDMI trained with CelebA-HQ at $\times64$ resolution (a). In (b), our model was trained on CelebA-HQ at $\times256$ using the hyper-transforming approach.
@@ 4.2 Generation
In Figure 5, we provide samples from LDMI trained on CelebA at $\times64$ and $\times256$, demonstrating its ability to generate high-quality and diverse images across different scales, and the qualitative superiority against the baselines. Additionally, in Table 1, we report FID scores of our samples for CelebA and ImageNet, highlighting the model’s performance in terms of image quality and reconstruction accuracy. Importantly, while GASP achieves a lower FID score on CelebA-HQ, it cannot perform reconstructions due to its adversarial design. In contrast, our LDMI supports both high-quality sampling and accurate reconstructions, as demonstrated in the following experiment.
@@ 4.3 Reconstruction
As for generation, our model enables seamless reconstruction at arbitrary resolutions. Figures 9 and 10 show results from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. To achieve this, we encode the observations, sample from the posterior distribution, and transform these latent samples into INRs using our HD decoder. The resulting INRs are then evaluated on grids of varying resolutions to produce reconstructions ranging from $\times 1/8$ to $\times 4$. Importantly, such flexible resolution control is not supported by methods like GASP, which cannot perform reconstructions due to their adversarial design.
To further validate our claims regarding cross-modality generalization, we evaluate LDMI on two additional, diverse datasets: ShapeNet Chairs and ERA5 climate data. As shown in Table 2, while Functa achieves better performance on Chairs—possibly due to being optimization-based—LDMI attains higher reconstruction quality than VAMoH (inference-based) across both datasets, outperforming it by 0.5% on Chairs and 5.6 dB on ERA5. The substantial performance gain on the ERA5 climate dataset highlights LDMI’s exceptional capability in representing complex, continuous spatio-temporal signals beyond standard visual data, further supporting our argument for a truly general-purpose INR-based generative framework.
Model Chairs (%) $\uparrow$ ERA5 (dB) $\uparrow$
Functa (Dupont et al., 2022a) 99.51 34.9
VAMoH (Koyuncu et al., 2023) 96.75 39.0
LDMI 97.25 44.6
Table 2: Reconstruction quality (accuracy in % and PSNR in dB) on ShapeNet Chairs and ERA5 climate data, demonstrating LDMI’s strong generalization capabilities across modalities. Note that GASP is omitted as it is not applicable to INR reconstruction tasks.
@@ Acknowledgements
This research was supported by the Villum Foundation through the Synergy project number 50091, by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (MLLS, grant NNF20OC0062606) and by the European Union (ERC-2021-STG, SAML, 101040177).
Ignacio Peis acknowledges support by Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429). Batuhan Koyuncu is supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education. The views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible.
@@ A.2 Diffusion Models
In DDPM, the forward diffusion process incrementally corrupts the observed variable ${\bm{z}}_{0}$ using a Markovian noise process with latent variables ${\bm{z}}_{1\mathrel{\mathop{\ordinarycolon}}T}$
$q({\bm{z}}_{1\mathrel{\mathop{\ordinarycolon}}T}|{\bm{z}}_{0})\mathrel{\mathop{\ordinarycolon}}=\prod_{t=1}^{T}q({\bm{z}}_{t}|{\bm{z}}_{t-1}),$ (18)
where $q({\bm{z}}_{t}|{\bm{z}}_{t-1})\mathrel{\mathop{\ordinarycolon}}=\mathcal{N}({\bm{z}}_{t};\sqrt{1-\beta_{t}}{\bm{z}}_{t-1},\beta_{t}\mathbf{I})$, and $\beta_{t}$ is a noise schedule controlling the variance at each time step.
@@
$p_{\theta}({\bm{z}}_{0\mathrel{\mathop{\ordinarycolon}}T})\mathrel{\mathop{\ordinarycolon}}=p({\bm{z}}_{T})\prod_{t=1}^{T}p_{\theta}({\bm{z}}_{t-1}|{\bm{z}}_{t})$ (19)
where $p_{\theta}({\bm{z}}_{t-1}|{\bm{z}}_{t})\mathrel{\mathop{\ordinarycolon}}=\mathcal{N}\left(\mu_{\theta}({\bm{z}}_{t},t),\bm{\sigma}_{t}^{2}\right)$ is defined as a Gaussian whose mean
@@ C.2 Samples at Multiple Resolutions
Figure 8 displays generations from LDMI trained on data across different modalities, rendered at scales ranging from $\times 0.125$ to $\times 4$. These results highlight the resolution-agnostic nature of our approach and its ability to produce coherent, high-quality outputs across a wide range of spatial resolutions.
@@
3D Chairs Refer to caption $\times 0.125$ Refer to caption $\times 0.25$ Refer to caption $\times 0.5$ Refer to caption $\times 1$ Refer to caption $\times 2$ Refer to caption $\times 4$
@@ C.3 Reconstructions at Multiple Resolutions
Figures 9 and 10 present reconstructions from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. Reconstructions are rendered at resolutions ranging from $\times 0.125$ to $\times 4$, showcasing the model’s ability to maintain spatial coherence and fine details even under extreme upsampling. Notably, such flexible reconstruction is not supported by baseline methods like GASP due to architectural limitations. Furthermore, datasets of this complexity are not even addressed by other reconstruction-capable baselines such as VAMoH (Koyuncu et al., 2023), Functa (Dupont et al., 2022b), or Spatial Functa (Bauer et al., 2023).
@@
$\times 0.125$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 0.25$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 0.5$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 1$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 2$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 4$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
$\times 0.125$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 0.25$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 0.5$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 1$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 2$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\times 4$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix D Experimental Details
resolution $\times 64 64×64$$\times 64$ $\times 256 256×256$$\times 256$ $46\times 90$ $32\times 32\times 32$
@@
latent_size $\times 16 16×16$$\times 16$ $\times 64 64×64$$\times 64$ $11\times 22$ $4\times 4$